Fermi National Laboratory
 

Accelerator Update for April 18, 2007 - April 20, 2007

 

Wednesday April 18

The day shift began with Operations monitoring the Recycler stash (272.9E10), with the Antiproton Source (Pbar) stacking (22.5E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

Operations put the Tevatron (TeV) into shot setup at 8:15 AM.

Operations established store 5358 at 10:21 AM with an initial luminosity of 214.19E30.

Operations reported at 11:03 AM that Pbar experts were conducting slipstacking studies, MI experts were conducting damper studies and about to begin MI-8 collimator studies, and Booster experts were about to begin quadrupole damper studies.

A Booster expert reported at 1:06 PM that the Notcher sounded like its charging resistor was arcing. He believed that the repairs would take about 90 minutes.

Operations established beam to Switchyard (SY) at 1:24 PM.

NuMI experts reported at 2:31 PM that they would take advantage of the Booster being down and begin repairs on a corrector magnet power supply (E:V118).

The Booster Notcher work was completed at 2:59 PM.

Operations reported at 4:17 PM that the Recycler had stashed antiprotons twice during the day shift.

Operations reported that the new Meson MTest experiment, T967, had received permission to begin taking beam. (T967 will use an electron beam and test the energy resolution, linearity, and shower size of a tungsten-scintillating fiber calorimeter segment.)

MI studies ended at 4:38 PM.

An MTest experimenter reported at 8:20 PM that they were through with the beam for the night.

At 10:59 PM, Operations reported problems with an Accumulator RF station (ARF4) during the last Recycler stash.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Thursday April 19

The midnight shift began with Operations monitoring store 5358 (55.38E30) and the stack (63.0E10), with Pbar stacking (43.8E10), and with NuMI and MiniBooNE taking beam.

At 12:16 AM, Operations contacted a Pbar expert about the ARF4 problems. He came in to investigate.

The Pbar expert reported at 2:15 AM that he had changed a blown fuse for ARF4.

At 2:55 AM, LRF2 tripped off on a driver overload. Operators reset the station.

At 3:17 AM, Pbar’s lithium lens tripped off on a power supply temperature fault. Operators reset the lens power supply and resumed stacking at 3:45 AM.

At 6:18 AM, Operations reported that ARF4 would not ramp.

Operation established beam to MTest at 6:56 AM.

At 7:07 AM, LRF2 tripped off. It reset and Linac resumed sending beam around 7:30 AM.

Operations reported at 7:59 AM that the Recycler had stashed twice during the midnight shift.

Operations terminated store 5358 at 8 AM.

Operations put the TeV into shot setup at 8:14 AM.

At 8:50 AM, Operations reported that ARF4 blew another fuse. Experts investigated.

Operations established store 5359 at 11:05 AM with an initial luminosity of 242.42E30.

At 11:47 AM, Linac expert began repairs on LRF2.

At 12:31 PM, a power supply expert began work on the NuMI corrector power supply E:V118.

LRF2 repairs were completed by 1:01 PM. Pbar, MiniBooNE, and MTest resumed taking beam.

At 1:36 PM, Operations reported a problem with beam falling out of MI at the start of ramp. Experts investigated and found a problem with their low level RF and damper systems. They rebooted both systems and resumed stable running at 2:18 PM.

Operations reported at 4 PM that the Recycler had stashed once during the day shift.

Recycler experts began an Electron Cooling study at 4:31 PM.

The MTest experiment reported at 6:01 PM that they were done with the beam for the night.

At 11:23 PM, Pbar’s lithium lens tripped off on a power supply high temperature.

At 11:28 PM, the TeV suffered a sector F4 quench. Operators investigated.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Friday April 20

The midnight shift began with the TeV in quench recovery, with Operations monitoring the stash (222.9E10), with Pbar stacking (5.2E10), and with NuMI and MiniBooNE taking beam.

At 12:27 AM, Operations reported more problems with LRF2. They called in an expert.

Operations prepared D0 and CDF for an access at 12:34 AM.

Operations reported at 12:39 AM that with directions from a Pbar expert operators switched over stashing responsibilities from ARF4 to ARF3.

Operations reported that LRF2 came back on around 1:12 AM. Pbar resumed stacking.

At 1:20 AM, LRF2 tripped off again.

CDF came out of access at 1:33 AM.

Operations reported at 2:34 AM that LRF2 was back up and running.

A TeV expert reported 3:18 AM that a CAMAC crate that held the sector B0 separators controls had its power supply short due to a bad control card. Experts replaced the card.

Operations put the TeV into a dry squeeze at 3:55 AM.

At 5 AM, Operations reported that they found a blown fuse on the sector B1 separator power supply controller.

At 5:09 AM, Operations loaded a 4x0 wet squeeze into the TeV.

At 6:12 AM, LRF2 tripped off. Operators reported that the station had no driver screen current. Experts were informed.

Operations reported at 8:06 AM that the Recycler had stashed once during the midnight shift.

The Plans for Friday and the Weekend

The plans for today are to get LRF2 fixed and establish a store. The plans for the weekend are to stack, stash, and store. The Run Coordinator may allow various machine studies over the weekend.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/20/2007    email Fermilab